OpenStudy (anonymous):

Solve the equation for y in terms of x: 3y + 2y^2 = -e^(-x) - e^(x) + 7

OpenStudy (dinamix):

its not differential equation ?

OpenStudy (anonymous):

Ill post the entire question and maybe that'll help

OpenStudy (anonymous):

Find the solution to the initial value problem in explicit form \[y' = \frac{e^{-x} - e^{x}}{3 + 4y}\]

OpenStudy (anonymous):

y(0) = 1

OpenStudy (anonymous):

@dan815 ?

OpenStudy (freckles):

y'*y doesn't equal y^2

OpenStudy (freckles):

\[(3+4y) dy=(e^{-x}-e^{x} )dx\] equation is seperable integrate both sides

OpenStudy (freckles):

oh wait are you saying you think you found the solution to that?

OpenStudy (anonymous):

yes i believe so and now i need it in explicit form

OpenStudy (anonymous):

I believe i have found the correct value for the constant c given the initial value as well

OpenStudy (freckles):

your solution is correct

OpenStudy (freckles):

don't believe you can find explicit form

OpenStudy (anonymous):

the answer in the back of the book has done it some how.

OpenStudy (freckles):

hmmm well you have a quadratic in terms of y

OpenStudy (freckles):

\[2y^2-3y+e^{-x}+e^{x}-7=0 \\ y=\frac{-(-3) \pm \sqrt{(-3)^2-4(2)(e^{-x}+e^{x}-7)}}{2(2)}\]

OpenStudy (freckles):

though the implicit form looks much better to me

OpenStudy (anonymous):

that looks like the answer, how'd you do that?!

OpenStudy (freckles):

it is a quadratic in terms of y

OpenStudy (freckles):

all i did was use the quadratic formula

OpenStudy (freckles):

a=2 b=-3 c=e^(-x)+e^x-7

OpenStudy (anonymous):

ohhhhh

OpenStudy (anonymous):

i gotcha now

OpenStudy (anonymous):

Thank you!!!!!!

OpenStudy (freckles):

np